the perimeter of a rectangle is 160ft.1/4 the length is the same as twice the width.Find the dimensions of the rectangle
Answer by mananth(16949)

1/2 the length = twice the width
multiply by 2
so L = 4w
Perimeter = 160=2(L+w)
160=2(4w+w)
160=10w
w= 16
L = 64
...
CHECK
16+16+64+64= 160